Circular No. 6165

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995L IN NGC 5157 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery with

the 0.9-m Schmidt telescope of a supernova located at α = 13^{h}27^{m}18^{s}.34, δ = +32^{o}01′52′′.7 (equinox 2000.0), which is 18′′.9 east and 2′′.0 north of the center of NGC 5157. A V plate taken on Mar. 26.05 UT shows the object at mag 20.0, while a B plate taken on Apr. 4.04 shows it at mag 19.5. Nothing is present at the position of SN 1995L on two films (limiting mag 22) obtained in 1994 Apr. A nearby star (mag ∼ 17) is at position end figures 09^{s}.04, 02′47′′.4. P. Garnavich, A. Riess, and P. Challis, Center for Astrophysics, report that a CCD image taken by R. Jansen with the 1.2-m telescope at Mt. Hopkins on Apr. 8.3 shows the supernova at V = 19.2 ± 0.1. A. V. Filippenko, A. J. Barth, and D. K. Stevens, University of California at Berkeley, report that preliminary inspection of an uncalibrated CCD spectrum (range 426–703 nm, resolution 0.7 nm) obtained on Apr. 22 UT with the 3-m Shane reflector at Lick Observatory reveals that the object is a supernova, most likely of type Ia several months past maximum brightness. The recession velocity of the host galaxy, measured from narrow emission lines in the nucleus, is 7260 km/s.

SUPERNOVA 1995F IN NGC 2726 A. Vagnozzi, G. Piermarini, and V. Russo, S. Lucia Observatory, Stron-

cone, Italy, report the following precise position for SN 1995F from a prediscovery CCD image obtained on Feb. 6.918 UT, when the object’s unfiltered magnitude was 14.3 (derived using GSC comparison stars): α = 9^{h}04^{m}57^{s}.40, δ = +59^{o}55′58′′.7 (equinox 2000.0).

V1413 AQUILAE F. Elizalde, Instituto Nacional de Pesquisas Espaciais (INPE), Sao

Jose dos Campos, Brazil; F. Jablonski, Laboratorio Nacional de Astrofisica (LNA), Itajuba; D. Cieslinski, INPE; and R. Baptista, Space Telescope Science Institute, report the following photometry obtained at the LNA’s Zeiss 0.6-m telescope on Apr. 2: V = 11.22(0.02), U −B = −0.23(0.03), B− V = +0.90(0.03), V −R = +0.78(0.02), R−I = +0.71(0.01). Simultaneous spectroscopy (range 590–625 nm, resolution 0.25 nm) obtained with the LNA 1.6-m telescope (+ Boller & Chivens spectrograph + CCD) confirms the description of Munari et al. (IAUC 6151), showing broad double-peaked Hα and He I lines with many weak lines of Fe II.

1995 April 22 Daniel W. E. Green